revenues are 1000; fixed manufacturing costs are $150; fixed non-manufacturing costs are $100; variable manufacturing costs are 250 and variable non-manufacturing costs are 200. the gross margin
Answer #1

Gross Profit = Sales - Cost of goods sold

Sales = 1000

Cost of goods sold = Fixed manufacturing cost + Variable manufacturing cost

Cost of goods sold = 150 + 250

Cost of goods sold = 400

Gross profit = 1000-400 i,.e 600

Gross margin = Gross profit / Sales *100

= 600/ 1000 *100 i.e 60%